OpenStudy (anonymous):

write the expression as a logarithm of a single quantity 3 ln x + 2 ln y - 4 ln z im guessing this one would be 1n(3x2y4z) (fan & award best answer)

OpenStudy (tkhunny):

Why are you guessing? \(3\ln(x) = \ln\left(x^{3}\right)\) Yours didn't look much like that. Give it another go.

OpenStudy (anonymous):

ln(x^3y^2 / z^4)

OpenStudy (tkhunny):

Well, that correct response was quite a bit differnt from your first response. What caused the change of heart?

OpenStudy (anonymous):

lolol ahaha,i didn't guess :)

OpenStudy (tkhunny):

Perfect. Not guessing is so much better than throwing darts. Good work. Let's make this a lesson learned and cut down on the guessing. Agreed?

OpenStudy (anonymous):

lol yes sir!